If square with a diagonal of 4√2 is equal to in area to a circle, that is the radius of the circle?
Answer by Jk22(389)

Diagonal = 4sqrt(2)=L*sqrt(2), L=square side length, hence L=4
Area(Square)=4*4=16
Area(Circle)=Pi*r^2, r is radius, still unknown.
hence equivalence leads to : Pi*r^2=16, and r=Sqrt(16/Pi)=2/Sqrt(Pi)